# Context is Key for Agent Security

### v2

@@ Abstract.
Judging the safety of an action, whether taken by a human or a system, must take into account the context in which the action takes place. For example, deleting an email from a user’s mailbox may or may not be appropriate depending on the email’s content, the user’s goals, or even available space. Today, existing systems (e.g., a smartphone ecosystem) that prevent harmful actions rely on manually-crafted policies or user confirmation for every context. With the upcoming deployment of generalist agents that support a multitude of tasks, we argue that we must rethink security designs to adapt to the scale of contexts and capabilities of these systems. As a first step, this paper explores contextual security in the domain of agents and proposes contextual security for agents (Conseca), a framework to generate just-in-time, contextual, and human-verifiable security policies.
copyright: none
@@ 1. Introduction
Context forms the basis for every action, whether taken by a human or a system: we can only disambiguate an action’s meaning via the context in which it exists. For example, in the context of an urgent deadline, scheduling over a lunch break might be appropriate, while the same action would be inappropriate for a casual sync. Researchers have integrated context to understand privacy in computer systems (Shvartzshnaider2019-nq; airgap; Kumar2020-lr; Nissenbaum2004PrivacyAC; Nissenbaum2019-kf; Apthorpe2018-dt; Barth2006-jb; Benthall2017-ur; Criado2015-bd; mireshghallah2023can; wijesekera2015android; grodzinsky2011privacy), which also manifests as appropriateness in AI agents (leibo2024theory). Importantly, in adversarial settings, analyzing surrounding context is necessary to differentiate harmful from benign actions.
However, approaches today for systems security depend on manually-specified contexts, or no context at all. Conventional systems security often uses policy enforcers and access controls (denningviews; griffithsauth; oracle-ac; zanzibar; iam; qapla; blockaid; predicated-ac; purpose-ac; finegrained-ac; oracle; oracle-ac; ms-sql; hippo; ifc-db). These can support fine-grained policies and capabilities, but rules are (manually) pre-written and policies static. Because policy or rule writers cannot account for every possible context, these rules can overrestrict actions in some contexts, overallow actions in others, and be challenging to write (reeder2007usability; johnson2010optimizing). Some systems’ reliance on e.g., manual user confirmation can also lead to user fatigue and overpermissioning (felt2012android; egelman2013choice; di2020ui; pixel; ootl; parkinson2019creeper). These limitations, while affecting any multi-purpose system like a filesystem or OS, have become increasingly important to address with the rise of systems like generalist (AI) agents (anthropic; aiassistants; reed2022generalistagent; frameworkagents; WooldridgeJennings), which may serve a far broader range of purposes and thus encounter far broader contexts.
@@
As a first step to explore how to scale security to match growing system capabilities, we focus on the domain of agents (AI and otherwise) and propose contextual security for agents (Conseca 111“Conseca” is the second-person singular present active imperative of the Latin verb conseco. Conseco can mean e.g., “to cut” or “to prune”.), a framework that moves us closer toward a world where every action has a contextually-appropriate, purpose-driven justification. Conseca demonstrates a possible way to integrate context at scale to achieve both better utility and better security than static security policies: Conseca generates just-in-time, contextual, and transparent security policies that specify which actions are not harmful for a given context and purpose (and restrict all other actions). Policy enforcement is deterministic, making enforcement impervious to attacks like prompt injections (pi1; owasp; willisonpi).
@@ 3.1. Trusted Context
While context is essential to judge an action’s appropriateness, adversaries can manipulate some of the context and thus also influence the policy generator model. Conseca relies on developers to specify what context to trust (typically context not directly under the control of arbitrary third parties); alternatively, Conseca could ask users to set coarse-grained trusted context boundaries. Trusting more context can allow Conseca to write a more accurate policy, but also increases the potential for compromise. For example, an email address could in principle encode malicious instructions (address formats typically allow long sequences (longurl)). However, a developer might consider email addresses from a monitored domain to be safe, enabling Conseca to write a better policy (e.g., restrict the agent to send emails to only “myteam@work.com” instead of any “*@work.com” address).
@@ 3.2. Contextual Security Policies
Developers can optionally ask users to approve a task’s policy prior to agent task execution, given the constraints and their human-readable rationales. Policies can also be logged and later audited by the user, the developer, or a trusted third party. Conseca relies on policy experts (perhaps automated) to ensure that the rationale matches the constraints.
@@ 3.4. Security Discussion
Conseca’s guarantees clearly depend on the strength of its security policies, which face two fundamental limitations. First, policies are only as good as the policy generator’s understanding of the user’s preferences, social norms, and actions required to fulfill the requested task. Second, the generator receives only limited, trusted contextual information (specified by the developer) in order to protect it against prompt injections. This restricts its ability to predict data-dependent actions (e.g., actions to allow when requested to “carry out the tasks requested in my manager’s email”). Because Conseca’s design uses an LLM, policies can improve by improving the model’s quality (tuning models for individual user preferences or on user’s feedback about generated policies). Conseca can also expand trusted context by e.g., asking the user or creating a sanitized language for action responses.
@@ 4.1. The Conseca Prototype
where a Policy maps an API call to constraints that include (i) whether the API call should ever be executed in this context, (ii) a boolean constraint over API call arguments such that the call can only execute when True; and (iii) a (human-readable) rationale for the choice of the prior two constraints. Policy generation invokes an LLM (Gemini 1.5 pro) with a prompt including the tool API documentation, example policies for in-context learning (icl), the user’s task, and the trusted context. For example, given:
@@
We (the “developers”) define trusted context as the users’ email categories and addresses, and a tree of the filesystem directory structure (file and directory names are trusted). Tool-agnostic context includes the user’s username, time, and date.
@@
Once Conseca sets the task’s policy, the agent invokes is_allowed when proposing a new command. Conseca checks whether the policy allows the API call at all, and, if so, whether each argument matches its regex constraint. Allowed commands go to execution. Conseca returns denied commands and their rationale to the planner, which the agent appends to the prompt to guide the planner LLM to choose an appropriate action. If commands continuously fail (up to 10), the agent returns “could not complete”.
@@ 5. Case Studies
Prior to running each task, we initialize the filesystem with 10 users, including an admin. Each user contains ¿10 files in both general and job-specific folders (e.g., Photos or Logs). Mailboxes start with emails from other users regarding work, family, etc.; some are categorized or include attachments like reports, invoices, and photos.
@@ 6. Related Work
Many agents’ use of LLMs leaves them vulnerable to jailbreaking (jailbreak) and prompt injections (pi1; owasp; ignore; willisonpi; indirectpi; ignore). Existing defenses based on e.g., heuristics (detect1; detect2; willisondetection; llamaguard) or anomaly detection (spotlighting; yidefense; willisondelimiters; instructionhierarchy) do not consider appropriateness of an action depending on the surrounding context. Conseca instead builds upon the idea of isolated agent models like SecGPT (secgpt) and f-secure LLM systems (wu2024systemleveldefenseindirectprompt), but in contrast, requires few changes to existing agents’ designs, explicitly aims to scale security to all contexts, and isolates policy generation rather than planning, thus providing full context for the planner.
Other work proposes a security analyser as part of agent design, which determines if a trace of agent actions violates a security policy composed of a set of pre-defined predicates (balunovic2024ai). Conseca proposes the idea of contextual policy generation independent of a specific policy language and enforcer, and can leverage their security analyser.
@@
Like prior work, Conseca uses models trained to understand behavioral norms (selfcorrection; secrets). Other works use these models for appropriateness (leibo2024theory) and in a contextual integrity (CI) framework (ci; Shvartzshnaider2019-nq; airgap; Kumar2020-lr; Nissenbaum2004PrivacyAC; Nissenbaum2019-kf; Apthorpe2018-dt; Barth2006-jb; Benthall2017-ur; Criado2015-bd; mireshghallah2023can; wijesekera2015android; grodzinsky2011privacy). Conseca follows in the spirit of CI, but enforces not just contextually appropriate data flows, but also contextually appropriate actions in order to prevent harm.
@@
Conseca helps take access control and capability restriction (denningviews; griffithsauth; oracle-ac; zanzibar; iam; qapla; blockaid; predicated-ac; purpose-ac; finegrained-ac; oracle; oracle-ac; ms-sql; hippo; ifc-db) to a finer granularity, which the rise of agent systems makes increasingly necessary.
Mobile device permissions today offer the best example of how a platform can capture user’s privacy and security preferences in a multitude of contexts. However, developers can abuse these permissions and users can over-grant permissions in order to use applications (permissionsandroid; androiddemystified). Agents serve even broader purposes than a mobile app, making it difficult to rely on users to manage permissions for every task; Conseca proposes a more scalable way to set permissions.
@@ 7. Discussion
Conseca only begins to touch upon the many directions for research in contextual security for agents. First, we can explore ways to improve contextual policies. While we hypothesize that users will experience far less confirmation fatigue with contextual policies than non-contextual ones, Conseca policies currently do not interact with users (e.g., asking users whether they want to override a Conseca-denied action, or giving users an “undo-log” that allows them to revert agent actions). User feedback can improve both the policy model’s understanding of preferences and norms, and also improve utility by allowing user-approved tasks to complete. To increase developers’ confidence in policies, we could perhaps automate policy verification using structured rationales and formally mapping them to constraints. Finally, sanitized output languages for tools and context may also increase the scope of trusted context and thus policy quality.
@@
Finally, can we generate contextual policies without LLMs at all? This might improve both policy reliability and efficiency: Conseca’s policy generator adds per-task LLM inference overheads for policy generation, which can take seconds depending on the size of the model. One alternative design, for example, might pre-generate and validate policies, and use them as a cache for similar tasks at agent runtime.
Exploring these directions for contextual security can help emerging agent systems leverage the power of context to choose and judge actions.

### v3

# Contextual Agent Security: A Policy for Every Purpose

@@ Abstract.
Judging an action’s safety requires knowledge of the context in which the action takes place. To human agents who act in various contexts, this may seem obvious: performing an action such as email deletion may or may not be appropriate depending on the email’s content, the goal (e.g., to erase sensitive emails or to clean up trash), and the type of email address (e.g., work or personal). Unlike people, computational systems have often had only limited agency in limited contexts. Thus, manually crafted policies and user confirmation (e.g., smartphone app permissions or network access control lists), while imperfect, have sufficed to restrict harmful actions. However, with the upcoming deployment of generalist agents that support a multitude of tasks (e.g., an automated personal assistant), we argue that we must rethink security designs to adapt to the scale of contexts and capabilities of these systems. As a first step, this paper explores contextual security in the domain of agents and proposes contextual agent security (Conseca), a framework to generate just-in-time, contextual, and human-verifiable security policies.
$c$s: Security and privacy Operating systems securityccs: Security and privacy Vulnerability managementccs: Security and privacy Access controlccs: Computing methodologies Artificial intelligencejournalyear: 2025copyright: ccconference: Workshop on Hot Topics in Operating Systems; May 14–16, 2025; Banff, AB, Canadabooktitle: Workshop on Hot Topics in Operating Systems (HOTOS ’25), May 14–16, 2025, Banff, AB, Canadadoi: 10.1145/3713082.3730378isbn: 979-8-4007-1475-7/2025/05
@@ 1. Introduction
Context forms the basis for every action, whether taken by a human or a system: we can disambiguate an action’s meaning only via the context in which it exists. For example, in the context of an urgent deadline, scheduling over a lunch break might be appropriate, while the same action would be inappropriate for a casual sync. Researchers have already integrated context to understand privacy in computer systems (Shvartzshnaider2019-nq; airgap; Kumar2020-lr; Nissenbaum2004PrivacyAC; Nissenbaum2019-kf; Apthorpe2018-dt; Barth2006-jb; Benthall2017-ur; Criado2015-bd; mireshghallah2023can; wijesekera2015android; grodzinsky2011privacy), which also manifests as appropriateness in AI agents (leibo2024theory). Importantly, in adversarial settings, analyzing surrounding context is necessary to differentiate harmful from benign actions.
However, approaches today for systems security depend on manually-specified contexts, or no context at all. Conventional security systems often use policy enforcers and access controls (denningviews; griffithsauth; oracle-ac; zanzibar; iam; qapla; blockaid; predicated-ac; purpose-ac; finegrained-ac; oracle; oracle-ac; ms-sql; hippo; ifc-db; NIST.SP.800-162). While these can support fine-grained policies and capabilities, their rules are (manually) prewritten and policies static. Because policy or rule writers cannot account for every possible context, these rules can overrestrict actions in some contexts, overallow actions in others, and be challenging to write (reeder2007usability; johnson2010optimizing). Some systems’ reliance on e.g., manual user confirmation can also lead to user fatigue and overpermissioning (felt2012android; egelman2013choice; di2020ui; pixel; ootl; parkinson2019creeper). These limitations, while affecting any multi-purpose system like a filesystem or OS, have become increasingly important to address with the rise of systems like generalist (AI) agents (anthropic; aiassistants; reed2022generalistagent; frameworkagents; WooldridgeJennings), which may serve a far broader range of purposes and thus encounter far broader contexts.
@@
As a first step to explore how to scale security to match growing system capabilities, we focus on the domain of agents (AI and otherwise) and propose contextual agent security (Conseca 111“Conseca” is the second-person singular present active imperative of the Latin verb conseco. Conseco can mean e.g., “to cut” or “to prune”.), a framework that moves us closer toward a world where every action has a contextually-appropriate, purpose-driven justification. Conseca demonstrates a possible way to integrate context at scale to achieve both better utility and better security than static security policies: Conseca generates just-in-time, contextual, and transparent security policies that specify which actions are not harmful for a given context and purpose (and restrict all other actions). Policy enforcement is deterministic, making enforcement impervious to attacks like prompt injections (pi1; owasp; willisonpi).
@@ 3.1. Trusted Context
While context is essential to judge an action’s appropriateness, adversaries can manipulate some of the context and thus also influence the policy generator model. Conseca relies on developers to specify what context to trust (typically context not susceptible to potential attackers); alternatively, Conseca could ask users to set coarse-grained trusted context boundaries. Trusting more context can allow Conseca to write a more accurate policy, but also increases the potential for compromise. For example, an email address could in principle encode malicious instructions (address formats typically allow long sequences (longurl)). However, a developer might consider email addresses from a monitored domain to be safe, enabling Conseca to write a better policy (e.g., restrict the agent to send emails to only “myteam@work.com” instead of any “*@work.com” address).
@@ 3.2. Contextual Security Policies
Developers can optionally ask users to approve a task’s policy prior to agent task execution, given the constraints and their human-readable rationales. Policies can also be logged and later audited by the user, the developer, or a trusted third party. Conseca relies on experts (perhaps automated) to ensure that the rationale matches the constraints.
@@ 3.4. Security Discussion
Conseca’s guarantees clearly depend on the strength of its security policies, which face two fundamental limitations. First, policies are only as good as the policy generator’s understanding of the user’s preferences, social norms, and actions required to fulfill the requested task. Second, the generator receives only limited, trusted contextual information (specified by the developer) in order to protect it against prompt injections. This restricts its ability to predict data-dependent actions (e.g., actions to allow when requested to “carry out the tasks requested in my manager’s email”). Because Conseca’s design uses an LLM, policies can improve by improving the model’s quality (tuning models for individual user preferences or on user’s feedback about generated policies). Conseca can also increase trusted context by e.g., asking the user or sanitizing action responses.
@@ 4.1. The Conseca Prototype
where a Policy maps an API call to constraints that include (i) whether the API call should ever be executed in this context, (ii) a boolean constraint over API call arguments such that the call can only execute when True; and (iii) a (human-readable) rationale for the choice of the prior two constraints. Policy generation invokes the LLM with a prompt including the tool API documentation, example policies for in-context learning (icl), the user’s task, and the trusted context. For example, given:
@@
We (the “developers”) define trusted context as the users’ email categories and addresses, and a tree of the filesystem directory structure (file and directory names are trusted). Tool-agnostic context includes the user’s username, time, and date. We also trust static context like the tool documentation provided in the prompt.
@@
Once Conseca sets the task’s policy, the agent invokes is_allowed when proposing a new command. Conseca checks whether the policy allows the API call at all, and, if so, whether each argument matches its regex constraint. Allowed commands go to execution. Conseca returns denied commands and their rationale to the planner, which the agent appends to the prompt to guide the planner LLM to choose an appropriate action. If commands continuously fail (up to 10 times), the agent returns “could not complete”.
@@ 5. Case Studies
Prior to running each task, we initialize the filesystem with 10 users, including an admin. Each user contains ¿10 files in each general or job-specific folder (e.g., Downloads, Photos, or Logs). Mailboxes start with emails from other users regarding work, family, etc.; some are categorized or include attachments like reports, invoices, and photos.
@@ 6. Related Work
Many agents’ use of LLMs leaves them vulnerable to jailbreaking (jailbreak) and prompt injections (pi1; owasp; ignore; willisonpi; indirectpi; ignore). Defenses that detect attacks (detect1; detect2; willisondetection; llamaguard), or prompt or train LLMs to distinguish untrusted from trusted inputs (spotlighting; yidefense; willisondelimiters; instructionhierarchy), remain susceptible to attacks introduced by untrusted context.
To address this susceptibility, other systems build upon a dual-LLM pattern (willisonisolation) that isolates the agent planner model from untrusted inputs (secgpt; wu2024systemleveldefenseindirectprompt). CaMeL (camel) strengthens this pattern by allowing the isolated planner to output a limited subset of Python, and utilizing a custom Python interpreter to track untrusted inputs and apply capability-based policies to deny or allow actions. Conseca also draws upon the idea of a quarantined, isolated model, but isolates policy generation rather than planning, thus providing full context for the planner and requiring few changes to existing agents’ planner designs. Conseca explicitly aims to scale security to all contexts by enabling dynamic and fine-grained security policies, and can help reduce user burden in these systems, which assume predefined security configurations or policies.
Other work proposes a security analyser as part of agent design, which determines if a trace of agent actions violates a security policy composed of a set of predefined predicates (balunovic2024ai). Conseca proposes the idea of contextual policy generation independent of a specific policy language and enforcer, and can leverage their security analyser.
@@
Like prior work, Conseca uses models trained to understand behavioral norms (selfcorrection; secrets). Other works use these models for appropriateness (leibo2024theory) and in a contextual integrity (CI) framework (ci; Shvartzshnaider2019-nq; airgap; Kumar2020-lr; Nissenbaum2004PrivacyAC; Nissenbaum2019-kf; Apthorpe2018-dt; Barth2006-jb; Benthall2017-ur; Criado2015-bd; mireshghallah2023can; wijesekera2015android; grodzinsky2011privacy). Conseca follows in the spirit of CI, but enforces contextually appropriate actions (in addition to contextually appropriate data flows) in order to prevent harm.
@@
Conseca helps enable finer-grained access controls and capability restrictions (denningviews; griffithsauth; oracle-ac; zanzibar; iam; qapla; blockaid; predicated-ac; purpose-ac; finegrained-ac; oracle; oracle-ac; ms-sql; hippo; ifc-db; apparmor; selinuxredhat), which the rise of agent systems makes increasingly necessary.
Mobile device permissions today offer the best example of how a platform can capture user’s privacy and security preferences in a multitude of contexts. However, developers can abuse these permissions and users might over-grant permissions in order to use applications (permissionsandroid; androiddemystified). Agents serve even broader purposes than a mobile app, making it difficult to rely on users to manage permissions for every task; Conseca proposes a more scalable way to set permissions.
Other systems for anomaly detection rely on manual behavior specification (saql) or learn a model of “good” behavior patterns (requiring labeled training data) (unicorn; anomalydetection). While more recent work has investigated LLMs for anomaly detection (su2024largelanguagemodelsforecasting), classic ML or behavioral models avoid the threat of prompt injection. A Conseca policy can be seen as a model of “good” behaviors for the current context, with deterministic enforcement to prevent prompt injections; in Conseca, the “model” is created by a pre-trained LLM rather than learned. Conseca targets agents that perform tasks expressed in natural language in unforeseen contexts, which naturally points towards the use of an LLM (which encodes natural language reasoning) for policy creation. These anomaly detection systems may operate better in orthogonal settings, where events do not include natural language (e.g., syscalls) and follow more constrained and predictable patterns.
@@ 7. Discussion
Conseca only begins to touch upon the many directions for research in contextual security for agents. First, we can explore ways to improve contextual policies. While we hypothesize that users will experience far less confirmation fatigue with contextual policies than non-contextual ones, Conseca policies currently do not interact with users (e.g., asking users whether they want to override a Conseca-denied action, or giving users an “undo-log” to audit agent actions or even revert them if possible). User feedback can improve both the policy model’s understanding of preferences and norms, and also improve utility by allowing user-approved tasks to complete. To increase developers’ confidence in policies, we could perhaps automate policy verification using structured rationales and formally mapping them to constraints. Finally, sanitized output languages for tools and context may also increase the scope of trusted context and thus policy quality.
@@
Use of LLMs also adds per-task overheads for policy generation, which can take seconds depending on the size of the model. LLM efficiency improvements, e.g., distilling a more capable model (xu2024survey), could reduce this cost, potentially trading off some quality. Alternatively, we could use caching techniques, storing pre-generated or dynamically created policies for common contexts.
Lastly, can we create contextual policies without LLMs at all? This could improve policy reliability and remove the problem of LLM hallucination (tonmoy2024comprehensive). The scope of agent contexts targeted by Conseca may not always apply: some agent systems may look more similar to existing limited-purpose applications, operate in restricted contexts, or not involve natural language, thus lending themselves to existing access control or anomaly detection techniques used today. In practice, even with the most general agents, developers would likely combine Conseca’s dynamic policies with expertly-written manual policies and/or user confirmation for high-risk scenarios (e.g., unlocking a door). During normal-risk operation, Conseca can provide more fine-grained (but less auditable) security policies while reducing both developer and user burden.
8. Conclusion
Conseca imagines an emerging world in which agents perform human-like tasks, serving different purposes and commanded by unstructured inputs. To match the scope of agents, our security systems must adapt to contexts beyond what can be manually or statically encoded. While not all deployed agents might achieve this scale of capabilities, more research into contextual agent security—safely leveraging the power of context to choose and judge actions—can help ensure that we are ready for such agents when they emerge.
9. Acknowledgements
We thank our anonymous reviewers for their feedback, as well as Borja De Balle Pigem, Christoph Kern, Jacint Szabo, Santiago Díaz, Andrew Baumann, Phil Levis, David Culler, Hank Levy, Marco Gruteser, Ilia Shumailov, Daniel Ramage, and others in Systems Research@Google, Google Research, Google DeepMind, and Google agent security efforts for their insights and discussion that led to this work.